OpenStudy (anonymous):

The sum of the squares of 2 consecutive negative integers is 41. What are the numbers? Which of the following equations is the result of using the factoring method to solve the problem? (n - 5)(n - 4) = 0 (n - 5)(n + 4) = 0 (n + 5)(n - 4) = 0 (n + 5)(n + 4) = 0

OpenStudy (anonymous):

Let the greater number be n. n^2 + (n - 1)^2 = 41. n^2 + n^2 + 1 - 2n = 41. 2n^2 - 2n - 40 = 0. n^2 - n - 20 = 0. Splitting the middle term, n^2 - 5n + 4n - 20 = 0 n(n - 5) + 4(n - 5) = 0 (n + 4)(n - 5) = 0. This is the answer to the second part of the question. Moving on, n + 4 = 0 or n - 5 = 0. So, n = - 4, 5. But, it is given that n is negative. So, n = -4 and n - 1 = -5. So, the numbers are - 5 and - 4.

OpenStudy (anonymous):

This assumes a well ordering for the integers.....